Find the magnetic flux through a 772 turn solenoid of length 33 cm, and radius 2 cm, carrying a current of 2 A.
Answer #1

magnetic field due to a solenoid is

B = mue0*ni/l = 4*3.14*10^-7*772*2/0.33 = 0.00587 T

flux = B*A = 0.00587*3.14*0.02^2 = 7.37*10^-6 wb